The average speed of a nitrogen molecule in air is about 6.70 ✕ 102 m/s, and its mass is about 4.68 ✕ 10-26 kg.

(a) If it takes 2.90 ✕ 10-13 s for a nitrogen molecule to hit a wall and rebound with the same speed but moving in an opposite direction (assumed to be the negative direction), what is the average acceleration of the molecule during this time interval?
m/s2

(b) What average force does the molecule exert on the wall?
N
Answer #1

Speed before collision (v1) = - 6.7 * 10^2 m/s

Speed after collision (v2) = + 6.7 * 10^2 m/s

Change in speed = v2 - v1 = 1.34 * 10^3 m/s

a)

Average acceleration = change in speed / time of collision

= (1.34 * 10^3 m/s) / (2.9 * 10^-13 s)

Average acceleration = 4.62 * 10^15 m/s^2

b)

Average force = mass * acceleration

= (4.68 * 10^-26 kg) * (4.62 * 10^15 m/s^2)

Average force = 2.16 * 10^(-10) N